Gucchi:

What is the length, in units, of the hypotenuse of a right triangle if each of the two legs is 2 units? 4 units 2 units Square root of 6 units Square root of 8 units

Gucchi:

Am I right with the last one?

AZ:

yes

AZ:

Good job!

pooja195:

A^2+b^2=c^2 2^2+2^2=?

Gucchi:

Answer: 8

Convert:

Okay for this one you would use the Pythagorean theorem and because 2^2 is 4 and 4+4 is 8 yes, I would believe that'd be correct! :)

pooja195:

Correct so you will have the square root of 8 here.

Gucchi:

Ok! Thank you.